40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 5 19:29:45 2012, Jenne, Update, General, Yarm aligned to IR incident beam 
    Reply  Thu Sep 6 02:51:47 2012, Koji, Update, General, Green beam roughtly aligned (Re: Yarm aligned to IR incident beam) green_y.png
       Reply  Thu Sep 6 10:57:27 2012, Jenne, Update, General, Forgot to turn green refl pd back on 
Message ID: 7346     Entry time: Wed Sep 5 19:29:45 2012     Reply to this: 7347
Author: Jenne 
Type: Update 
Category: General 
Subject: Yarm aligned to IR incident beam 

[EricQ, Jenne, brains of other people]

Checked at ETMY that the pointing hadn't changed a whole lot.  Jamie and Koji pointed out that we were half falling off of the IPANG QPD.  Adjusted PZT2 sliders so that we were again centered on IPANG's QPD.  Before we close up, we'll want to put the sliders back to their nominal positions, and use the knobs to hit IPANG, but this is equivalent and fine for now.  The tip tilts don't seem to have moved much overnight, since the beam drift on both IPANG and ETMY was fixed simultaneously with PZT2 (recall, IPANG picked off before tip tilts exist in the beam path).  This left us hitting the center of ETMY.  We moved ETMY sliders to make the reflected beam hit the center of ITMY (same spot position as transmitted beam from BS).  Then moved ITMY to get prompt reflection to hit same spot on ETMY as original primary beam from BS.  Checked at ITMY that we didn't need to move ETMY anymore.  (Actually, I forget how many iterations we did, but in the end, all of the reflections that we can find are co-located on the test masses.)

Next up:

Align BS so we're hitting the center of ETMX

Tap / readjust ITMX OSEM which is at 0.3 to get it back to the center of its range

Align ITMX to lock MICH

Check no clipping on POX / POY optics, no clipping around BS

Check PRM, SRM alignment (what exactly do we want to do here? Try to lock PRMI / SRMI?)

Get IPPOS out of vac

Fix clipped ITMY / SRM oplev

Install 'black' glass beam dumps - forward-going POP beam, 2 places in BS chamber (check old elog from Jenne/Yuta for the places).

Get green spots co-located with IR spots on ETMs, ITMs, check path of leakage through the arms, make sure both greens get out to PSL table

ELOG V3.1.3-